Find two consecutive numbers such that 3/4 of the first number is 5 more than 1/2 the second number
Answer by ankor@dixie-net.com(6693)

Two consecutive numbers: x, (x+1)
:
Find two consecutive numbers such that 3/4 of the first number is 5 more than 1/2 the second number.
x = (x+1) + 5
;
Get rid of the denominators, multiply all terms by 4, results
3x = 2(x+1) + 4(5)
:
3x = 2x + 2 + 20
:
3x - 2x = 22
:
x = 22 & 23 are the numbers
;
:
Check solution in the statement:
"two consecutive numbers such that 3/4 of the first number is 5 more than 1/2 the second number."
22 = (23) + 5
16.5 = 11.5 + 5